A hospital receives 20% of its COVID-19 vaccine shipments from Japan and the
remainder of its shipments from neighbouring countries. Each shipment contains a very large number of vaccine vials. For Japan’s shipments, 10% of the vials are ineffective. For the neighbouring countries, 2% of the vials are ineffective.
The hospital tests 30 randomly selected vials from a shipment and finds that one is ineffective.
What is the probability that the shipment came from Japan ?
1
Expert's answer
2020-05-19T16:18:56-0400

H1={from Japan}; p(H1)=0.2

H2={from neighbouring countries}; p(H2)=0.8

A={1 is ineffective from 30 randomly selected}

"p(H_1/A)=\frac{p(A/H_1)p(H_1)}{p(A/H_1)p(H_1)+p(A/H_2)p(H_2)}"

"p(A\/H_1)=C_{30}^1(0.1)(0.9)^{29}=0.1413"

"p(A\/H_2)=C_{30}^1(0.02)(0.98)^{29}=0.334"

"p(H_1/A)=\frac{0.1413*0.2}{0.1413*0.2+0.8*0.334}=0.096"

Answer: 0.096.